From 9 mathematicians and 6 physicists a committee of 4 mathematicians and 3 physicists is to be formed.What is the probability that commitee has
a) 3 particular mathematicians b)one particular physicist

Answer by stanbon(48568)

From 9 mathematicians and 6 physicists a committee of 4 mathematicians and 3 physicists is to be formed.What is the probability that commitee has
a) 3 particular mathematicians b)one particular physicist
---------------
Mathematicians:
I'm going to assume the selector knows which 9 are mathematicians and
which 6 are physicists. If he/she doesn't the problem is quite different.
---------------
Pick the special three then pick one more to get four:
P(pick the three) = 1/9C3 = 1/64
P(pick one more) = 6/6 = 1
===================
Physicists:
Pick the special one then pick two others:
P(pick the special one) = 1/6
P(pick 2 others) = (5/5)(4/4)=1
============================
Prob(the combined event) = 1/64 * 1/6 = 1/504 = 0.00198...
===============================
Let me know if this is the answer you have and/or whether
my assumption is wrong. If you have another solution I
would like to see it.
-----------------------------
Cheers,
Stan H.